# Knowledge-aware Parsimony Learning: A Perspective from Relational Graphs

### v2

@@ Abstract
[Abstract] The scaling law, which involves the brute-force expansion of training datasets and learnable parameters, has become a prevalent strategy for developing more robust learning models. However, due to bottlenecks in data, computation, and trust, the sustainability of the scaling law is a serious concern for the future of deep learning. In this paper, we address this issue by developing next-generation models in a parsimonious manner (i.e., achieving greater potential with simpler models). The key is to drive models using domain-specific knowledge, such as symbols, logic, and formulas, instead of relying on the scaling law. This approach allows us to build a framework that uses this knowledge as “building blocks” to achieve parsimony in model design, training, and interpretation. Empirical results show that our methods surpass those that typically follow the scaling law. We also demonstrate the application of our framework in AI for science, specifically in the problem of drug-drug interaction prediction. We hope our research can foster more diverse technical roadmaps in the era of foundation models.
@@ 1 Introduction
The learning techniques have progressed from manual feature engineering to shallow models, then to deep networks, and now to foundation models, achieving great success in the field of computer vision, natural language understanding and speech processing. Specifically, large language models, like ChatGPT (Ouyang et al., 2022), as representatives of foundation model, has shown strong performance in versatile learning, which can adopted in many different tasks. The belief is that larger models can be more expressive, thus are likely to generalize better given sufficient training data (Jordan and Mitchell, 2015; Donoho, 2017). This gives birth to the current roadmap, i.e., achieving stronger performance by aggressively scaling up the size of data and models, which is also observed as scaling law (Kaplan et al., 2020).
@@
However, such a roadmap potentially leads to serious problems (as shown in the left of Figure 1): Data bottleneck—the scaling law relies on vast amounts of high-quality data, yet all available online corpora are projected to be exhausted by 2028 (Villalobos et al., 2024); Computational bottleneck—the exponentially growing number of parameters demands substantial high-performance computing power, yet the pace of hardware development struggles to keep up with this rapid increase (Wolters et al., 2024); and Trust bottleneck—the scaling law follows the data driven path, disregarding internal logical relationships, which leads to opaque reasoning processes and severe hallucination issues (Duan et al., 2024). These indicate that the current roadmap is not sustainable, and motivates us to ask:
@@
Intuitively, humans accumulate knowledge in the form of symbols, concepts, rules, and principles, which allows us to learn a wide range of subjects or skills, apply them across various tasks, quickly adapt to new tasks with few or no demonstrations, and uncover the underlying reasons behind different phenomena. Inspired by (Lake et al., 2015; Cranmer et al., 2020; Lake and Baroni, 2023), which shows that neural networks can achieve human-like systematicity with smaller models and achieve performance comparable to that of larger models, we propose knowledge driven methods that emphasize guiding machine learning models with domain-specific knowledge, using it as “building blocks” to achieve parsimony in learning. This approach allows us to tackle the bottlenecks of scaling law, focusing instead on efficient model design, training, and interpretation. We demonstrate the effectiveness of this framework in AI for science, particularly in addressing the problem of drug-drug interaction prediction. Empirical results show that our methods can outperform those limited by the scaling law.
@@ 2 Research Landscape
Our key innovation lies in enabling parsimony learning through knowledge-aware approaches. Specifically, knowledge, i.e., symbolic logic and physical laws, is ubiquitous in real-world scenarios and coexists with data. As illustrated on the left side of Figure 2, the diverse colors found in nature can be derived from three primary colors (i.e., red, blue, and green). Building on this, we propose the concept of the “duality of knowledge and data”, which posits that data is both numerical and knowledge-based. Based on this concept, we design a framework capable of learning from both data and knowledge simultaneously, thereby effectively achieving parsimony in learning. As shown on the right side of Figure 2, the framework first identifies symbolic logic and physical laws at the knowledge level, then performs combinatorial generalization of this knowledge at the data level. By treating knowledge as “building blocks” and focusing on its learning and generalization, the framework leverages simple knowledge to solve complex problems.
To efficiently address these bottlenecks, we introduce domain-specific knowledge into our framework to achieve greater potential with smaller models. Specifically, we implement model parsimony, leveraging simple architectures to overcome computational bottleneck. Additionally, training parsimony is employed to reduce training costs, tackling the data bottleneck. To address the trust bottleneck, interpretive parsimony is used to identify key evidence. Finally, we apply this proposed framework to the field of AI for science, demonstrating its significant potential.
@@
  • Parsimony on Model. The goal of architectural parsimony is to utilize simple architecture to achieve the comparable performance of complex models. As introduced in (Lake and Baroni, 2023), simple neural networks can achieve human-like systematicity when optimized for compositional skills, offering the potential for smaller models to achieve performance comparable to that of larger ones. Following this intuition, we propose AutoBLM (Zhang et al., 2022) and ColdNAS (Wu et al., 2023), which utilize prior knowledge to streamline architectures. By achieving this, we can maintain high performance with simpler structures, thereby efficiently addressing computational bottlenecks.
@@
  • Parsimony on Training. The purpose of training parsimony is to reduce the amount of training data. Inspired by (Ilharco et al., 2023), which shows that task-related knowledge help to determine parameters by simple arithmetic operations without training, we propose new approaches for learning parsimony (Wang et al., 2021; Yao et al., 2024; Wu et al., 2024). These approaches leverage knowledge to guide the fine-tuning process on specific tasks under few-shot circumstances. By achieving this, we can efficiently utilize knowledge to guide the optimization of related tasks with limited training data, thereby efficiently addressing data bottlenecks.
@@
  • Parsimony on Interpretation. The parsimony on interpretation is to identify important evidence in face to massive connections on graphs. Our primary approach is to encapsulate logical rules within graphs as supporting evidence, utilizing the logical interpretation of knowledge to elucidate the reasoning processes of models (Shi et al., 2023; Zhang and Yao, 2022). By achieving this, we can efficiently interpret the model’s result with subgraph, thereby efficiently addressing trust bottlenecks.
  • Potential in AI for Science. In above, we have elaborate how parsimony in model, training and interpretation can be achieved individually following a knowledge-driven way. We will combine these idea and further show how they can help AI for science. Specifically, we will take drug-drug interaction as an application, which predicts interactions between emerging and existing drugs. By extracting path-based subgraphs and learning subgraph representation, we can efficiently reduce the data and computational requirements and identify evidence with subgraphs.
@@ 3.1 Automated Bi-linear Scoring Function Design
Table 1: MRR performance comparison of AutoBLM and neural models (Interstellar and CompGCN).
@@ 4 Parsimony on Training
In AI-assisted scientific research, particularly in molecules and biomedicine, the scarcity of labeled data presents significant challenges. Motivated by (Ilharco et al., 2023), which utilizes the symbolic structure between parameters obtained from different tasks for downstream tasks prediction, we propose approaches to learning parsimony by leveraging task-related knowledge to guide the fine-tuning on downstream tasks. In molecules properties prediction, we have developed meta-learning learning techniques that enforce parsimony on learning. These approaches ensure that parameters can be efficiently adapted in relation to functional groups, optimizing their use and enhancing model adaptability.
@@
4.1 Property-Aware Relation Networks
4.2 Parameter-Efficient GNN Adapter
Table 3 shows the prediction performance. Our analysis reveals that methods utilizing pretrained graph-based molecular encoders generally outperform those with encoders learned from scratch. This underscores the effectiveness of pretrained encoders in capturing rich, generic molecular information, subsequently providing superior molecular embeddings. In further evaluations, meta-learning methods that learn relational graphs—specifically EGNN, PAR, and PACIA—demonstrate enhanced performance. Notably, PACIA consistently achieves the highest ROC-AUC scores, followed closely by PAR. In summary, PACIA establishes itself as the new state-of-the-art for predicting molecular properties. It not only delivers superior predictive accuracy but also features significantly faster adaptation speed and requires fewer adaptive parameters.
@@ 5.1 Interpreting with Subgraph Learning
Figure 10: (a) EmerGNN learns pairwise representations of drugs by extracting the paths between drug pairs, propagating information from one drug to the other, and incorporating the relevant biomedical concepts on the paths. (b) Visualization of the structure learned by EmerGNN. DB006204 (Tapentadol) is an existing drug, and DB00757 (Dolasetron) is an emerging drug. (c) Performance comparison on DDI prediction tasks. The best results are bolded and the second-best underlined.
@@ 6 Potential in Drug Development
The general idea of EmerGNN is illustrated in Figure 10. Emerging drugs typically have limited interactions with existing drugs. To address this issue, EmerGNN utilizes biomedical network and extracts subgraph $\mathcal{G}_{u,v}^{L}$ from it to connect emerging drug $u$ with existing drug $v$. Drawing inspiration from the subgraph learning methods (Zhang and Yao, 2022; Zhu et al., 2021), EmerGNN proposes a flow-based GNN $g(\mathcal{G}_{u,v}^{L};\bm{\theta})$ with an attention mechanism to encode pair-wise subgraph representations for drug pairs. These subgraph representations are then used to directly predict the interaction of the drug pair. This approach enables a more nuanced understanding of interactions involving emerging drugs, incorporating knowledge from the interconnected network of biomedical entities and their relationships, and providing interpretable insights.
The results are shown in Table 10(c), from the results, we find that the proposed EmerGNN outperforms CSMDDI (Liu et al., 2022) that directly uses drug features to predict, HINDDI (Tanvir et al., 2021) that counts meta-paths from biomedical network for prediction, KG-DDI (Karim et al., 2019) that learns drug embeddings, and DeepLGF (Ren et al., 2022) that models with GNN in both benchmarks and both settings. Furthermore, we show a case study with the visualized subgraph in Figure 10(b) by selecting important paths according to the attention values. The path connecting two drugs through the binding protein Gene::1565 (CYP2D6), which is a P450 enzyme that plays a key role in drug metabolism, for example, shows a way to interpret the inference process with supporting evidence. From the empirical comparison and case studies, we conclude that subgraph learning can not only achieve improvements in representation learning, but also interpret the inference process with supporting evidence.
@@ 7 Future Works
  • Theory: Knowledge quantization. One key area for further research is the development of a theoretical foundation for knowledge quantization. Current machine learning frameworks heavily depend on large-scale data and complex models; however, incorporating mechanisms to distill knowledge into more compact and efficient forms can result in models that are both more efficient and parsimonious. Future work will focus on quantifying how much domain-specific knowledge is necessary to drive model performance while reducing the need for large datasets. This will involve formalizing the concept of “knowledge as a resource” and developing algorithms that can optimize the utilization of knowledge while maintaining performance across various tasks. A theoretical framework for knowledge quantization will pave the way for more efficient integration of symbolic logic, physical laws, and other forms of structured knowledge into learning models.
  • Methods: Integrating with LLMs. As large language models (LLMs) continue to dominate AI research, finding efficient ways to integrate them with specialized knowledge and smaller models is crucial. Our future work will focus on creating methods that allow seamless integration of LLMs with domain-specific frameworks. Instead of scaling up models indefinitely, the focus will be on using LLMs as modular components that can be fine-tuned or adapted based on task-specific knowledge. This approach will involve the development of techniques for aligning the outputs of LLMs with symbolic and logical reasoning systems, enhancing the ability of these models to perform in specialized fields such as bioinformatics or physics. Furthermore, exploring ways to reduce the computational footprint of LLMs while maintaining their versatility will be a key area of focus.
  • Applications: AI for Science. The application of AI for scientific discovery remains one of the most promising areas for further exploration. Future research will focus on leveraging the proposed parsimonious learning framework to tackle more complex challenges in drug discovery, molecular property prediction, and other areas within AI for science. In drug-drug interaction prediction, the intricate nature of molecular interactions typically demands significant computational resources. By applying parsimony to the model, neural networks can be simplified, enabling efficient predictions while reducing computational load. In protein structure prediction, the scarcity of labeled data often makes training challenging. Parsimony in training leverages domain knowledge, thereby reducing the reliance on large datasets and enabling accurate predictions even with limited data. In molecular property prediction, understanding the reasoning behind model predictions is crucial for advancing scientific research. Parsimony in interpretation allows for the extraction of key insights from complex data, improving both the transparency of the model and the trustworthiness of its predictions.

### v3

# Beyond Scaleup: Knowledge-aware Parsimony Learning from Deep Networks

@@ Abstract
[Abstract] The brute-force scaleup of training datasets, learnable parameters and computation power, has become a prevalent strategy for developing more robust learning models. However, due to bottlenecks in data, computation, and trust, the sustainability of this strategy is a serious concern. In this paper, we attempt to address this issue in a parsimonious manner (i.e., achieving greater potential with simpler models). The key is to drive models using domain-specific knowledge, such as symbols, logic, and formulas, instead of purely relying on scaleup. This approach allows us to build a framework that uses this knowledge as “building blocks” to achieve parsimony in model design, training, and interpretation. Empirical results show that our methods surpass those that typically follow the scaling law. We also demonstrate our framework in AI for science, specifically in the problem of drug-drug interaction prediction. We hope our research can foster more diverse technical roadmaps in the era of foundation models.
@@ 1 Introduction
The learning techniques have progressed from manual feature engineering to shallow models, then to deep networks, and now to foundation models, achieving great success in the field of computer vision, natural language understanding and speech processing. Specifically, large language models, like ChatGPT (Ouyang et al., 2022), as representatives of foundation model, has shown strong performance in versatile learning, which can adopted in many different tasks. The belief is that larger models can be more expressive, thus are likely to generalize better given sufficient training data (Jordan and Mitchell, 2015; Donoho, 2017). This gives birth to the current roadmap, i.e., achieving stronger performance by aggressively scaleup the size of data and models, which is also observed as scaling law (Kaplan et al., 2020).
@@
However, such a roadmap potentially leads to serious problems (as shown in the left of Figure 1): Data bottleneck—the scaling law relies on vast amounts of high-quality data, yet all available online corpora are projected to be exhausted by 2028 (Villalobos et al., 2024); Computational bottleneck—the exponentially growing number of parameters demands substantial high-performance computing power, yet the pace of hardware development struggles to keep up with this rapid increase (Wolters et al., 2024); and Trust bottleneck—the scaling law follows the data driven path, disregarding internal logical relationships, which leads to opaque reasoning processes and severe hallucination issues (Duan et al., 2024). These indicate that the current roadmap is not sustainable, and motivates us to ask: where is the way if scaleup fails?
@@
Intuitively, humans accumulate knowledge in the form of symbols, concepts, rules, and principles, which allows us to learn a wide range of subjects or skills, apply them across various tasks, quickly adapt to new tasks with few or no demonstrations, and uncover the underlying reasons behind different phenomena. Inspired by (Lake et al., 2015; Cranmer et al., 2020; Lake and Baroni, 2023), which shows that neural networks can achieve human-like systematicity with smaller models and achieve performance comparable to that of larger models, we propose knowledge driven methods that emphasize guiding machine learning models with domain-specific knowledge, using it as “building blocks” to achieve parsimony in learning. This approach allows us to tackle the bottlenecks of scaleup, focusing instead on efficient model design, training, and interpretation. We demonstrate the effectiveness of this framework in AI for science, particularly in addressing the problem of drug-drug interaction prediction. Empirical results show that our methods can outperform those limited by the scaling law.
@@ 2 Research Landscape
Our key innovation lies in enabling parsimony learning through knowledge-aware approaches. Specifically, knowledge, i.e., symbolic logic and physical laws, is ubiquitous in real-world scenarios and coexists with data. As illustrated on the left side of Figure 2, the diverse colors found in nature can be derived from three primary colors (i.e., red, blue, and green). Building on this, we propose the concept of the “duality of knowledge and data”, which posits that data is both numerical and knowledge-based. We subsequently design a framework capable of learning from both data and knowledge simultaneously, thereby effectively achieving parsimony in learning. As shown on the right side of Figure 2, the framework first identifies symbolic logic and physical laws at the knowledge level, then performs combinatorial generalization of this knowledge at the data level. By treating knowledge as “building blocks” and focusing on its learning and generalization, the framework leverages simple knowledge to solve complex problems.
Here, we incorporate domain-specific knowledge into our framework and propose parsimony on model to overcome computational bottlenecks, parsimony on training to mitigate data bottlenecks, and parsimony on interpretation to address trust bottlenecks. We then explore the potential of our approach in AI for Science, aiming to tackle all these bottlenecks within a unified framework. As shown in Table 1, we illustrate the knowledge space, function space, and their relationships in each part.
Table 1: Illustration of knowledge space, function space, and the relationship in each method.
@@
  • Parsimony on Model. The goal of architectural parsimony is to utilize simple architecture to achieve the comparable performance of complex models. As introduced in (Lake and Baroni, 2023), simple neural networks can achieve human-like systematicity when optimized for compositional skills, offering the potential for smaller models to achieve performance comparable to that of larger ones. We use bi-level optimization to simultaneously search architecture in knowledge space and obtain representation in functional space to achieve parsimony on model. Following this intuition, we propose AutoBLM (Zhang et al., 2022) and ColdNAS (Wu et al., 2023), which utilize prior knowledge to streamline architectures. By achieving this, we can maintain high performance with simpler structures, thereby efficiently addressing computational bottlenecks.
@@
  • Parsimony on Training. The purpose of training parsimony is to reduce the number of parameters needs to be updated during the training. Inspired by (Thrun and Pratt, 1998; Rastogi, 2022; Ilharco et al., 2023), which shows that task-related knowledge help to determine parameters by simple arithmetic operations without training, we propose new approaches for learning parsimony (Wang et al., 2021; Yao et al., 2024; Wu et al., 2024). Specifically, we use neural networks to learn semantic relationships across different tasks in knowledge space and learn to update parameters in representation learning in function space adaptively. These approaches leverage knowledge to guide the fine-tuning process on specific tasks under few-shot circumstances. By achieving this, we can efficiently utilize knowledge to guide the optimization of related tasks with limited training data, thereby efficiently addressing data bottlenecks.
@@
  • Parsimony on Interpretation. The parsimony on interpretation is to identify important evidence in face to massive connections on graphs. Our primary approach is to encapsulate logical rules within graphs as supporting evidence, utilizing the logical interpretation of knowledge to elucidate the reasoning processes of models (Shi et al., 2023; Zhang and Yao, 2022). We use knowledge space to constraint the learning process in the functional space and then use performance obtained from functional space to identify keys in the knowledge space. By achieving this, we can efficiently interpret the model’s result with subgraph, thereby efficiently addressing trust bottlenecks.
  • Potential in AI for Science. To efficiently co-evolve the knowledge and functional spaces to address the bottlenecks of data, computation, and interoperability, we propose a unified framework specifically designed for drug development. To enhance the interpretability of drug-drug interactions, we employ neural networks to extract biomedical relationships between subgraphs and drugs within the knowledge space. These relationships are then utilized for drug-pair interaction prediction within the functional space. Additionally, by leveraging simple architectures (subgraphs) identified in the knowledge space and neural networks for learning molecular representations, we achieve superior performance compared to traditional methods. Specifically, we take drug-drug interaction as an application, which predicts interactions between emerging and existing drugs. By extracting path-based subgraphs and learning subgraph representation, we can efficiently reduce the data and computational requirements and identify evidence with subgraphs.
@@ 3.1 Automated Bi-linear Scoring Function Design
Table 2: MRR performance comparison of AutoBLM and other neural models.
@@ 4 Parsimony on Training
In AI-assisted scientific research, particularly in molecules and biomedicine, the scarcity of labeled data presents significant challenges. Motivated by (Palatucci et al., 2009; Rastogi, 2022; Ilharco et al., 2023), which utilizes the symbolic structure between parameters obtained from different tasks for downstream tasks prediction, we propose approaches to learning parsimony by leveraging task-related knowledge to guide the fine-tuning on downstream tasks. In molecules properties prediction, we have developed meta-learning learning techniques that enforce parsimony on learning. These approaches ensure that parameters can be efficiently adapted in relation to functional groups, optimizing their use and enhancing model adaptability.
4.1 Property-Aware Relation Networks
PACIA PAR
# Total para. 3.28M 2.31M
# Adaptive para. 3.00K 0.38M
Table 4: Total and adaptive parameters of PACIA and PAR in few-shot learning.
4.2 Parameter-Efficient GNN Adapter
@@
Table 5 shows the prediction performance. Our analysis reveals that methods utilizing pretrained graph-based molecular encoders generally outperform those with encoders learned from scratch. This underscores the effectiveness of pretrained encoders in capturing rich, generic molecular information, subsequently providing superior molecular embeddings. In further evaluations, meta-learning methods that learn relational graphs—specifically EGNN, PAR, and PACIA—demonstrate enhanced performance. Notably, PACIA consistently achieves the highest ROC-AUC scores, followed closely by PAR. In summary, PACIA establishes itself as the new state-of-the-art for predicting molecular properties. Additionally, as shown in Table 4, PACIA can obtain better performance due to the reduction of adaptive parameters, which also leads to better generalization and alleviates the risk of overfitting to a few shots.
@@ 5.1 Interpreting with Subgraph Learning
Figure 10: (a) EmerGNN learns pairwise representations of drugs by extracting the paths between drug pairs, propagating information from one drug to the other, and incorporating the relevant biomedical concepts on the paths. (b) Performance and data amount comparison between traditional, transformer, and LLM-based methods.
Figure 11: (a) Visualization of the structure learned by EmerGNN. DB006204 (Tapentadol) is an existing drug, and DB00757 (Dolasetron) is an emerging drug. (b) Performance comparison on DDI prediction tasks. The best results are bolded and the second-best underlined. (c) The number of parameters comparison between EmerGNN, Graph Transformer, and LLM-based methods.
@@ 6 Potential in Drug Development
The general idea of EmerGNN is illustrated in Figure 10(a). Emerging drugs typically have limited interactions with existing drugs. To address this issue, EmerGNN utilizes biomedical network and extracts subgraph $\mathcal{G}_{u,v}^{L}$ from it to connect emerging drug $u$ with existing drug $v$. Drawing inspiration from the subgraph learning methods (Zhang and Yao, 2022; Zhu et al., 2021), EmerGNN proposes a flow-based GNN $g(\mathcal{G}_{u,v}^{L};\bm{\theta})$ with an attention mechanism to encode pair-wise subgraph representations for drug pairs. These subgraph representations are then used to directly predict the interaction of the drug pair. This approach enables a more nuanced understanding of interactions involving emerging drugs, incorporating knowledge from the interconnected network of biomedical entities and their relationships, and providing interpretable insights. We also compare the performance and data requirements of traditional, transformer-based, and LLM-based methods, with the results shown in Figure 10(b). By learning pairwise representations of drugs, EmerGNN achieves superior performance while using less data compared to other methods.
The performance comparison on DDI prediction tasks is shown in Figure 11(b). From the results, we find that the proposed EmerGNN outperforms CSMDDI (Liu et al., 2022) that directly uses drug features to predict, HINDDI (Tanvir et al., 2021) that counts meta-paths from biomedical network for prediction, KG-DDI (Karim et al., 2019) that learns drug embeddings, and DeepLGF (Ren et al., 2022) that models with GNN in both benchmarks and both settings, TIGER (Su et al., 2024) that leverages the Transformer architecture to effectively exploit the structure of heterogeneous graph. TextDDI (Zhu et al., 2023) uses a language model-based DDI predictor and a reinforcement learning (RL)-based information selector. Furthermore, we show a case study with the visualized subgraph in Figure 11(a) by selecting important paths according to the attention values. The path connecting two drugs through the binding protein Gene::1565 (CYP2D6), which is a P450 enzyme that plays a key role in drug metabolism, for example, shows a way to interpret the inference process with supporting evidence. From the empirical comparison and case studies, we conclude that subgraph learning can not only achieve improvements in representation learning, but also interpret the inference process with supporting evidence. Additionally, we compare the number of parameters across different methods, with the results shown in Figure 11(c). By incorporating biomedical relationships between subgraphs and drugs within the knowledge space, we effectively reduce the number of training parameters while maintaining superior performance.
@@ 7 Future Works
  • Theory. Current machine learning relies on large-scale data and complex models, but distilling knowledge into compact forms can make models more efficient. Future work will quantify the necessary domain-specific knowledge to enhance model performance while reducing large datasets. This involves formalizing “knowledge as a resource” and creating algorithms to optimize knowledge use while maintaining performance across tasks. A theoretical framework for knowledge quantization will enable better integration of symbolic logic, physical laws, and structured knowledge into learning models.
  • Methods. As LLMs dominate AI research, integrating them with specialized knowledge and smaller models is crucial. Future work will focus on methods for seamless integration of LLMs with domain-specific frameworks. Instead of indefinitely scaleup models, the focus will be on using LLMs as modular components that can be fine-tuned based on task-specific knowledge. This approach involves developing techniques to align LLM outputs with symbolic and logical reasoning systems, enhancing performance in specialized fields like bioinformatics or physics. Additionally, reducing the computational footprint of LLMs while maintaining their versatility will be a key focus.
  • Applications. AI’s application in scientific discovery is highly promising. Future research will use parsimonious learning to address complex challenges in drug discovery, molecular property prediction, and other scientific areas. In drug-drug interaction prediction, simplifying neural networks can reduce computational demands. For protein structure prediction, leveraging domain knowledge can improve accuracy despite limited data. In molecular property prediction, parsimony enhances model transparency and trustworthiness by extracting key insights from complex data.
Through endeavor in above directions, we aim to develop a unified and general methodology where knowledge representations and functional spaces co-evolve synergistically. This approach involves designing frameworks that leverage task-specific domain knowledge to simultaneously reduce model complexity, minimize training parameters, and enhance interpretability. The goal is for knowledge to refine the functional space — such as model architectures and representations — while the learned functions provide feedback to enrich and expand the knowledge space. This unified methodology promises to create more adaptable, interpretable, and efficient learning systems.